samantha ran on a fitness track fo 20 minutes and then walked for 10 minutes. her running rate was 250 ft per minute faster than her walking rate. if she ran and walked a total of 15,500 ft, how fast did she run?
Answer by richwmiller(17219)

r=w+250,
20r+10*w=15500
20(w+250)r+10*w=15500
20w+5000+10w=15500
30w=10500
3w=1050
w=350 f/m
r=600 f/m